tickets to the fashion show are 3.50 for adults and 1.50 for students the total sales were 767.50 for a total of 301 tickets that were sold how many students and adults attended the fashion show?
Answer by mananth(16946)

x= number of students
y= number of adults

1 x + 1 y = 301 .............1

1.50 x + 3.50 y = 767.50 .............2
Eliminate y
multiply (1)by -3.50
Multiply (2) by 1.00
-3.50 x -3.50 y = -1053.50
1.50 x + 1.00 y = 767.50
Add the two equations
-2.00 x = -286.00
/ -2.00
x = 143.00
plug value of x in (1)
1.00 x + 1.00 y = 301.00
143.00 + y = 301.00
y = 301.00 -143
y = 158.00
y = 158.00 1
x= 143 number of students
y= 158 number of adults
m.ananth@hotmail.ca